Show and explain each step on how to solve the logarithmic equation.

1) log2(x)=4
2) lnx=2.7
3) ln(x)=3
4) 6.27=e^x

show all work pleas







Answer by ankor@dixie-net.com(22740)

Show and explain each step on how to solve the logarithmic equation.
1) log2(x)=4
write the exponent equiv of given log
x = 2^4
x = 16
:
2) lnx=2.7
Write the exponent equiv using e
x = e^2.7
use your calc here
x = 14.8797
:
3) ln(x)=3
same here
x = e^3
x = 20.0855
:
4) 6.27=e^x
Find the ln of both sides
ln(6.27) = ln(e^x)
log equiv of exponents
ln(6.27) = x*ln(e)
ln of e = 1, therefore
ln(6.27) = x
use your calc, find ln of 6.27
x = 1.835776